OpenStudy (anonymous):

if ln k = 0.7, calculate the value of the following expression: \[\ln \sqrt[3]{k}/10 + \ln(10k^{2})\]

OpenStudy (anonymous):

use these: 1)ln(a/b) = lna -lnb 2)ln(a^b) = b lna 3)ln(ab) = lna + lnb try now..

OpenStudy (campbell_st):

using the log runs above \[\frac{3}{2} \log (k) - \log(10) + \log(10) + 2 \log(k)\] hopefullt this will help you find the answer.

OpenStudy (anonymous):

32log(k)−log(10)+log(10)+2log(k) =3/2 log (0,7) - log (10) + log (10) + 2 log (0,7) =3/2 -0,15 -1 +1 + 1,4 =4,25/2 @champbell_st @Aryang plz check this ans is it rite?